Fulbright Corp. uses the periodic inventory system. During its first year of operations, Fulbright made the following purchases (listed in chronological order of acquisition):

  • 41 units at $104 per unit
  • 74 units at $88 per unit
  • 171 units at $53 per unit


Sales for the year totaled 265 units, leaving 21 units on hand at the end of the year.

Ending inventory using the average cost method is: (Do not round unit cost calculation. Round your final answer to the nearest whole dollar amount.)

Multiple Choice

  • A $2,184.

  • B $1,163.

  • C $1,457.

  • D $1,113.
Answer #1

Answer : C = $ 1457.

Average cost method - Periodic inventory Method.

>> Cost of goods available for sale = ( 41 * 104 ) + ( 74 * 88 ) + ( 171 * 53 ) = $ 19839.

>>Total goods available = 41 + 74 + 171 = 286.

>>Average cost = ( $ 19839 / 286 ) = $ 69.3671

>>Closing inventory = 21 * $ 69.3671 = $ 1457